# ImprovNet: Generating Controllable Musical Improvisations with Iterative Corruption Refinement

### v1

@@ Abstract
Deep learning has enabled remarkable advances in style transfer across various domains, offering new possibilities for creative content generation. However, in the realm of symbolic music, generating controllable and expressive performance-level style transfers for complete musical works remains challenging due to limited datasets, especially for genres such as jazz, and the lack of unified models that can handle multiple music generation tasks. This paper presents ImprovNet, a transformer-based architecture that generates expressive and controllable musical improvisations through a self-supervised corruption-refinement training strategy. ImprovNet unifies multiple capabilities within a single model: it can perform cross-genre and intra-genre improvisations, harmonize melodies with genre-specific styles, and execute short prompt continuation and infilling tasks. The model’s iterative generation framework allows users to control the degree of style transfer and structural similarity to the original composition. Objective and subjective evaluations demonstrate ImprovNet’s effectiveness in generating musically coherent improvisations while maintaining structural relationships with the original pieces. The model outperforms Anticipatory Music Transformer in short continuation and infilling tasks and successfully achieves recognizable genre conversion, with 79% of participants correctly identifying jazz-style improvisations. Our code and demo page can be found at https://github.com/keshavbhandari/improvnet.
@@ VII-A Generalizability and Link to Diffusion Models
The corruption-refinement framework used in ImprovNet highlights its potential applicability to broader domains, including large language models (LLM) or other unstructured data tasks, as a generalizable approach for iterative improvement. This method shares conceptual parallels with diffusion models, where corruptions can be interpreted as analogous to the noise applied during the diffusion process. While training diffusion models, noise is sampled from a predefined probabilistic distribution, while in ImprovNet, corruptions are sampled from a uniform distribution. The number of refinement passes in ImprovNet mirrors the iterative denoising steps of diffusion models. However, ImprovNet provides an additional advantage: users can specify the corruption function, allowing for greater controllability over the “noise” introduced, a level of customization not typically explored in diffusion models. A key distinction lies in the generation process: ImprovNet employs an autoregressive approach, where notes are generated in an ordered sequence, and each segment is iteratively refined, leveraging its temporal context. In contrast, diffusion models generally predict an entire sequence or representation at once, guided by the iterative noise removal process.

### v3

# ImprovNet - Generating Controllable Musical Improvisations with Iterative Corruption Refinement

@@ Abstract
Despite deep learning’s remarkable advances in style transfer across various domains, generating controllable performance-level musical style transfer for complete symbolically represented musical works remains a challenging area of research. Much of this is owed to limited datasets, especially for genres such as jazz, and the lack of unified models that can handle multiple music generation tasks. This paper presents ImprovNet, a transformer-based architecture that generates expressive and controllable musical improvisations through a self-supervised corruption-refinement training strategy. The improvisational style transfer is aimed at making meaningful modifications to one or more musical elements - melody, harmony or rhythm of the original composition with respect to the target genre. ImprovNet unifies multiple capabilities within a single model: it can perform cross-genre and intra-genre improvisations, harmonize melodies with genre-specific styles, and execute short prompt continuation and infilling tasks. The model’s iterative generation framework allows users to control the degree of style transfer and structural similarity to the original composition. Objective and subjective evaluations demonstrate ImprovNet’s effectiveness in generating musically coherent improvisations while maintaining structural relationships with the original pieces. The model outperforms Anticipatory Music Transformer in short continuation and infilling tasks and successfully achieves recognizable genre conversion, with 79% of participants correctly identifying jazz-style improvisations of classical pieces. Our code and demo page can be found at https://github.com/keshavbhandari/improvnet.
@@ VII-A Generalizability and Link to Diffusion Models
The corruption-refinement framework used in ImprovNet highlights its potential applicability to broader domains, including large language models (LLM) or other unstructured data tasks, as a generalizable approach for iterative improvement. A single pass of our method shares conceptual parallels with inpainting in Stable Diffusion [^4] https://stable-diffusion-art.com/inpainting_basics/, where corruptions can be interpreted as analogous to the mask applied during the diffusion process. However, the multiple refinement passes in ImprovNet resembles the process of repeating an entire masked diffusion model multiple times (equal to the number of refinement passes), akin to performing several full denoising cycles. However, ImprovNet provides an additional advantage: users can specify the corruption function, allowing for greater controllability over the “noise” introduced, a level of customization not typically explored in diffusion models. A key distinction lies in the generation process: ImprovNet employs an autoregressive approach, where notes are generated in an ordered sequence, and each segment is iteratively refined, leveraging its temporal context. In contrast, diffusion models generally predict an entire sequence or representation at once, guided by the iterative noise removal process.